# Multi-Task Dense Prediction via Mixture of Low-Rank Experts

### v1

@@ 2.1 Dense Multi-Task Learning
In computer vision, multi-task learning (MTL) for dense prediction tasks has been widely studied. The previous methods can be divided into two categories, including optimization-based methods and architecture-based methods [48]. Optimization methods [28, 8, 18, 59, 9] facilitate MTL by utilizing different strategies to balance the influence of each task. Architecture-based methods aim to design a unified deep network for MTL. They can be further classified into two categories, encoder-focused methods, and decoder-focused methods. (normal-(((i)normal-))) Encoder-focused methods [40, 34, 3, 19] design multi-task backbones to extract features adapted to different tasks. Typical methods include cross-stitch networks [36], neural discriminative dimensionality reduction networks [17], multi-task attention networks [31], branched networks [47], and mixture-of-expert networks [15, 10]. (normal-(((ii)normal-))) Decoder-focused methods [50, 49, 4, 52, 53, 54, 58, 57, 60] share the same backbone and design delicate heads to extract task-specific features for each task and cross-task relationships. The advantage of decoder-focused methods lies in that they can benefit from powerful off-the-shelf vision backbones, such as DINOv2 [37]. Our method also falls into the decoder-focused category and studies how to produce task-specific features with the MoE technique.
@@ 3.2 Preliminaries: Mixture of Experts
[…]
@@ 3.3 Mixture of Low-Rank Experts
Task-sharing low-rank expert path. We take inspiration from LoRA [22] and adopt low-rank convolution which is a low-rank format of a vanilla convolution. Each task-sharing expert network shares a similar structure consisting of a $3\times 3$ convolution and a $1\times 1$ convolution. The weights and bias of all task-sharing expert networks can be formulated as $\left\{\mbox{{W}}^{n}_{lreb},\mbox{{b}}^{n}_{lreb},\mbox{{W}}^{n}_{lrea},\mbox{{b}}^{n}_{lrea}|n\in[1,...,N]\right\}$, where $\mbox{{W}}^{n}_{lreb}\in\mathbb{R}^{3\times 3\times C\times r_{n}},\mbox{{b}}^{n}_{lreb}\in\mathbb{R}^{r_{n}},\mbox{{W}}_{lrea}\in\mathbb{R}^{1\times 1\times r_{n}\times C}$ , and $\mbox{{b}}^{n}_{lrea}\in\mathbb{R}^{C}$ ($r_{n}\ll C$). $r_{n}$ denotes the rank for the $n$-th expert network. In our method, $r_{n}$ for different expert networks is different, aiming to improve the diversity of parameters and features. For each task, the task-specific router network $f^{t}_{sr}(\cdot)$ learns gate values for these experts and activate the top-$k$ experts according to the gate values. The output features of all the activated experts are summed up and then sent to a BatchNorm layer to generate task-specific features. The BatchNorm layer contains four parameters, including the accumulated channel-wise mean $\mathbf{\mu}\in\mathbb{R}^{C}$, the accumulated channel-wise standard deviation $\mathbf{\sigma}\in\mathbb{R}^{C}$, the scaling factor $\gamma\in\mathbb{R}^{C}$ and the $\beta\in\mathbb{R}^{C}$, respectively.
@@ A.1 MLoRE Module Stacking
In our method, we stack two MLoRE modules after each selected backbone layer. In the first MLoRE module, the lightweight task-specific $1\times 1$ convolutions are utilized to project the backbone feature to different task features. In the second MLoRE module, as the task-specific features have been split already, we utilize $1\times 1$ convolutions directly to deal with task-specific features. Moreover, as the MLoRE is a linear module, we add a task-specific non-linear block between the two MLoRE modules to introduce the non-linearity into our decoder. Each non-linear block is composed of a GELU-BatchNorm-Linear structure.

### v2

@@ 2.1 Dense Multi-Task Learning
In computer vision, multi-task learning (MTL) for dense prediction tasks has been widely studied. The previous methods can be divided into two categories, including optimization-based methods and architecture-based methods [48]. Optimization methods [28, 8, 18, 59, 9] facilitate MTL by utilizing different strategies to balance the influence of each task. Architecture-based methods aim to design a unified deep network for MTL. They can be further classified into two categories, encoder-focused methods, and decoder-focused methods. ((((i)))) Encoder-focused methods [40, 34, 3, 19] design multi-task backbones to extract features adapted to different tasks. Typical methods include cross-stitch networks [36], neural discriminative dimensionality reduction networks [17], multi-task attention networks [31], branched networks [47], and mixture-of-expert networks [15, 10]. ((((ii)))) Decoder-focused methods [50, 49, 4, 52, 53, 54, 58, 57, 60] share the same backbone and design delicate heads to extract task-specific features for each task and cross-task relationships. The advantage of decoder-focused methods lies in that they can benefit from powerful off-the-shelf vision backbones, such as DINOv2 [37]. Our method also falls into the decoder-focused category and studies how to produce task-specific features with the MoE technique.
@@ 3.2 Preliminaries: Mixture of Experts
For M3ViT [15] and Mod-Squad [10], they deactivate some experts according to the corresponding gate values for one inference and select the top-$k$ experts. The task-specific features are used to make predictions for each task.
@@ 3.3 Mixture of Low-Rank Experts
Task-sharing low-rank expert path. We take inspiration from LoRA [22] and adopt low-rank convolution which is a low-rank format of a vanilla convolution. Each task-sharing expert network shares a similar structure consisting of a $3\times 3$ convolution and a $1\times 1$ convolution. The weights and bias of all task-sharing expert networks can be formulated as $\left\{\mbox{{W}}^{n}_{lreb},\mbox{{b}}^{n}_{lreb},\mbox{{W}}^{n}_{lrea},\mbox{{b}}^{n}_{lrea}|n\in[1,...,N]\right\}$, where $\mbox{{W}}^{n}_{lreb}\in\mathbb{R}^{3\times 3\times C\times r_{n}},\mbox{{b}}^{n}_{lreb}\in\mathbb{R}^{r_{n}},\mbox{{W}}^{n}_{lrea}\in\mathbb{R}^{1\times 1\times r_{n}\times C}$ , and $\mbox{{b}}^{n}_{lrea}\in\mathbb{R}^{C}$ ($r_{n}\ll C$). $r_{n}$ denotes the rank for the $n$-th expert network. In our method, $r_{n}$ for different expert networks is different, aiming to improve the diversity of parameters and features. For each task, the task-specific router network $f^{t}_{sr}(\cdot)$ learns gate values for these experts and activate the top-$k$ experts according to the gate values. The output features of all the activated experts are summed up and then sent to a BatchNorm layer to generate task-specific features. The BatchNorm layer contains four parameters, including the accumulated channel-wise mean $\mathbf{\mu}\in\mathbb{R}^{C}$, the accumulated channel-wise standard deviation $\mathbf{\sigma}\in\mathbb{R}^{C}$, the scaling factor $\gamma\in\mathbb{R}^{C}$ and the $\beta\in\mathbb{R}^{C}$, respectively.
@@ A.1 MLoRE Module Stacking
In our method, we stack two MLoRE modules after each selected backbone layer. In the first MLoRE module, the lightweight task-specific $1\times 1$ convolutions are utilized to project the backbone feature to different task features. In the second MLoRE module, as the task-specific features have been split already, we utilize $1\times 1$ convolutions directly to deal with task-specific features. Moreover, as the MLoRE is a linear module, to introduce the non-linearity into the decoder, we insert the task-specific non-linear block between the two MLoRE modules and after the last MLoRE module. Each non-linear block is composed of a BatchNorm-GELU-Linear structure. When fusing the features from different scales, we generate a pixel-wise mask from the concatenated features and leverage it to weight the feature from different scales.
@@ A.2 MoE Optimization
At last, we also enable the router to predict one extra scaling value to weight the feature from the task-specific low-rank expert path. The scaling value doesn’t get involved in top-$k$ selecting process.